A mother is 36 years old and her son is 6 years. When will the product if their ages be 451.

Answer by ikleyn(52803)

.

Find the decomposition of the number 451 into the product of two natural numbers with the difference  36-6 = 30 between them.


You can do it MENTALLY:  451 = 41*11.


Answer.  In 5 years.


Alternatively,  you can solve the quadratic equation

        (36+x)*(6+x) = 451

and to get the same answer.